# Evidence of isospin-symmetry violation in high-energy collisions of atomic nuclei – Supplementary Information

### v3

@@ 7.2 Violation in the charm sector
Similar to kaons, exact charge-symmetry invariance predicts an equal production of charged mesons, $D^{+}$ and $D^{-}$, and neutral ones, $D^{0}$ and $\bar{D}^{0}$. However, the production of charged $D^{\pm}$ mesons in $h+A$ interactions is suppressed relatively to $D^{0}$ and $\bar{D}^{0}$ [Frixione:1997ma]. A similar suppression observed in nucleus–nucleus collisions at RHIC and at the LHC. For example, in Au+Au collisions at $\sqrt{s_{\rm NN}}=200\,\text{GeV}$ (10–40% centrality), the STAR measurement finds midrapidity cross sections $d\sigma_{D^{0}}/dy\simeq 41\,\mu\text{b}$ and $d\sigma_{D^{\pm}}/dy\simeq 18\,\mu\text{b}$, giving a yield ratio [Vanek:2021axf, Xie:2017nal]
@@
At the LHC, the ALICE measurements of prompt $D^{0}$ and $D^{+}$ production in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02\,\text{TeV}$ show the ratio $\langle D^{+}\rangle/\langle D^{0}\rangle\approx 0.4$$0.5$ [ALICE:2018lyv], and is compatible (within uncertainties) with the same ratio in $p+p$ interactions [ALICE:2021rxa].
Following the same reasoning as in the kaon case, this is a manifestation of charge-symmetry-breaking. However, in the charm case, there is a simple explanation of the breaking. The charged vector resonance $D^{*}(2010)^{+}$ decays into $D^{+}\pi^{0}$ and $D^{0}\pi^{+}$ with branching ratios of $30.7\pm 0.5\%$ and $67.7\pm 0.5\%$, respectively. Their ratio is close to 0.5, as expected from isospin-symmetry invariance. On the other hand, the resonance $D^{*}(2007)^{0}$ decays into $D^{0}\pi^{0}$ with branching ratio of $64.7\pm 0.9\%$, but the decay into its charged partner, $D^{-}\pi^{+}$, is forbidden because it is below the threshold. Exact isospin symmetry would imply a $1:2$ ratio between $D^{0}\pi^{0}$ and $D^{-}\pi^{+}$ decays. The effect is qualitatively similar to $\phi(1020)$ resonance decay in the case of kaons. Quantitatively, the impact of $D^{*}$ decays on the charged-to-neutral $D$ ratio is much larger, since the resonance $D^{*}(2007)^{0}$ does not decay at all into $D^{+}$. Moreover, this resonance has a significant ($35.3\pm 0.9\%$) branching ratio to $D^{0}\gamma$, which also increases neutral $D$ meson production w.r.t. charged $D$ mesons. Similar considerations hold for the charge-conjugated states $D^{*}(2010)^{+}$ and $\bar{D}^{*}(2007)^{0}$. Summarising, neutral $D$ mesons are significantly favoured (by a factor of about two) over charged ones.
@@ Acknowledgments
The authors thank Wojciech Broniowski, Katarzyna Grebieszkow, Stanislaw Mrowczynski, Owe Philippsen, Adam Para, Rob Pisarski, Krishna Rajagopal, Jan Steinheimer, Leonardo Tinti, Ludwik Turko and Volodymyr Vovchenko for fruitful discussions and comments. A special thank you goes to Herbert Stroebele for suggestions, corrections, and remarks. This work is partially supported by the Polish National Science Centre (NCN) grants 2018/30/A/ST2/00226 and 2019/33/B/ST2/00613 and the Alexander von Humboldt Foundation.

### v4

@@ 7.2 Violation in the charm sector
Similar to kaons, exact charge-symmetry invariance predicts an equal production of charged mesons, $D^{+}$ and $D^{-}$, and neutral ones, $D^{0}$ and $\bar{D}^{0}$. However, the production of charged $D^{\pm}$ mesons in $h+A$ interactions is suppressed relative to $D^{0}$ and $\bar{D}^{0}$ [21]. A similar suppression was observed in nucleus–nucleus collisions at RHIC and at the LHC. For example, in Au+Au collisions at $\sqrt{s_{\rm NN}}=200\,\text{GeV}$ (10–40% centrality), the STAR measurement finds mid-rapidity cross sections $d\sigma_{D^{0}}/dy\simeq 41\,\mu\text{b}$ and $d\sigma_{D^{\pm}}/dy\simeq 18\,\mu\text{b}$, giving a yield ratio [53, 60]
$R_{D}=\frac{\langle D^{+}\rangle+\langle D^{-}\rangle}{\langle D^{0}\rangle+\langle\bar{D}^{0}\rangle}\approx 0.5\text{ .}$ (27)
At the LHC, the ALICE measurements of prompt $D^{0}$ and $D^{+}$ production in Pb–Pb collisions at $\sqrt{s_{\rm NN}}=5.02\,\text{TeV}$ show the ratio $\langle D^{+}\rangle/\langle D^{0}\rangle\approx 0.4$$0.5$ [3], and is compatible (within uncertainties) with the same ratio in $p+p$ interactions [4].
Following the same reasoning as in the kaon case, this is a manifestation of charge-symmetry-breaking. However, in the charm case, there is a simple explanation for the breaking. The charged vector resonance $D^{*}(2010)^{+}$ decays into $D^{+}\pi^{0}$ and $D^{0}\pi^{+}$ with branching ratios of $30.7\pm 0.5\%$ and $67.7\pm 0.5\%$, respectively. Their ratio is close to 0.5, as expected from isospin-symmetry invariance. On the other hand, the resonance $D^{*}(2007)^{0}$ decays into $D^{0}\pi^{0}$ with branching ratio of $64.7\pm 0.9\%$, but the decay into its charged partner, $D^{-}\pi^{+}$, is forbidden because it is below the threshold. Exact isospin symmetry would imply a $1:2$ ratio between $D^{0}\pi^{0}$ and $D^{-}\pi^{+}$ decays. The effect is qualitatively similar to $\phi(1020)$ resonance decay in the case of kaons. Quantitatively, the impact of $D^{*}$ decays on the charged-to-neutral $D$ ratio is much larger, since the resonance $D^{*}(2007)^{0}$ does not decay at all into $D^{+}$. Moreover, this resonance has a significant ($35.3\pm 0.9\%$) branching ratio to $D^{0}\gamma$, which also increases neutral $D$ meson production w.r.t. charged $D$ mesons. Similar considerations hold for the charge-conjugated states $D^{*}(2010)^{+}$ and $\bar{D}^{*}(2007)^{0}$. Summarising, neutral $D$ mesons are significantly favoured (by a factor of about two) over charged ones.
In Fig. 4 we show the HRG result for $R_{D}$. At high energies, the model yields $R_{D}\simeq 0.45$, in good agreement with the experimental value $R_{D}\simeq 0.5$. This agreement indicates that the HRG framework provides a reliable baseline description for charmed mesons. The fact that the same framework fails to reproduce the charged-to-neutral kaon ratio highlights the non-trivial nature of the $R_{K}$ puzzle.
Figure 4: Ratio $R_{D}$ within HRG for $Q/B=0.5$ and including resonances up to $D^{*}(2010)$. Note, using $Q/B=0.4$ leads to negligible numerical differences.
8 Closing remarks
In this work, triggered by the novel experimental findings of Ref. [6] and discussion therein, we have investigated the ratio $R_{K}$ of charged-to-neutral kaon yields in nucleus-nucleus collisions at high energies, both within the general theoretical QCD framework and within the Hadron Resonance Gas model.
In the limit of exact $ud$-flavour symmetry, the $R_{K}$ ratio is unity if the initial state of nuclei has an equal number of protons and neutrons ($Q/B=1/2$). Here, we have didactically clarified the origin of the result, presented a formal proof, and discussed differences to the Shmushkevich method.
The known charge symmetry violation effects included in HRG yield $R_{K}\approx 1.03$ at high collision energies. Many experimental results correspond to collisions of heavy nuclei with more neutrons than protons ($Q/B\simeq 0.4$). This imbalance reduces $R_{K}$ significantly only at low collision energies, where the ratio is lower than one. We thus conclude that the known charge-symmetry-violating effects cannot explain experimental results on nucleus-nucleus collisions, for which $R_{K}$ is mostly within the range 1.1-1.2.
The evidence for unexpectedly large CSV in the kaon sector is well established. It is, however, important that new experimental tests take place in the future. Measurements of the charged-to-neutral kaon ratio in collisions of an equal number of protons and neutrons would reduce interpretation uncertainties. In this respect, the NA61/SHINE collaboration initiated an investigation of the $R_{K}$ ratio in oxygen-oxygen scattering. In addition, measurements of $R_{K}$ in deuterium-deuterium interactions, the lightest nuclei with $I=0$ and $Q/B=1/2$ would also be important.
Another interesting system concerns pion-nucleon scattering, in particular $\pi^{+}+C$ and $\pi^{-}+C$. Previous data on $\pi^{-}+C$ have shown that $R_{K}\simeq 1.2$ [5], in line with nucleus-nucleus scattering. Charge symmetry implies that the average $R_{K}$ obtained by combining both measurements should be close to unity. On the other hand, the coalescence approach for hadronic production [29] shows that both systems should lead to the same enhanced $R_{K}$ with $R_{K}^{\pi^{+}C}=R_{K}^{\pi^{-}C}\simeq 1.2>1$.
On the theoretical side, different directions can be taken. Concerning the HRG approach, one may argue that isospin-breaking contributions have not yet been included. This scenario seems, however, improbable: the higher the mass of a resonance, the smaller its multiplicity. Moreover, such resonances should exhibit a large, unexplained isospin-symmetry-breaking decay to account for the experimentally measured $R_{K}$.
@@ 8 Closing remarks
It is expected that, in the near future, both experimental and theoretical advances will help clarify the origin of the isospin kaon anomaly.
@@ Acknowledgments
The authors thank Wojciech Broniowski, Katarzyna Grebieszkow, Stanislaw Mrowczynski, Owe Philippsen, Adam Para, Rob Pisarski, Krishna Rajagopal, Jan Steinheimer, Leonardo Tinti, Ludwik Turko, and Volodymyr Vovchenko for fruitful discussions and comments. A special thank you goes to Herbert Stroebele for suggestions, corrections, and remarks. This work is partially supported by the Polish National Science Centre (NCN) grants 2018/30/A/ST2/00226 and 2019/33/B/ST2/00613 and the Alexander von Humboldt Foundation.